Physics college

A cylindrical rod of uniform density is located with its centre at the origin and its axis along the x axis. It rotates about its centre in the xy plane, making one revolution every 0.02 s. The rod has a radius of 0.02 m, length of 0.2 m, and mass of 5 kg. What is the rotational kinetic energy of the rod?
Answer #1

KErot = 1/2 * I * w2

I = MR2/4 + ML2/12


I = 5*0.022/4 + 5*0.22/12 = 0.017166


w = 2π/T  = 2π/0.02


Therefore, KErot = 1/2 * 0.017166* 4π2/0.0004 = 847.141 J

answered by: Saanvi Singh